all the information are provided2) A debt of $32 000 is repaid by payments of $2950 made at the end of every six months. Interest is 8.28% compounded quarter
Answer #1

Question a:

Quarterly interest rate = 8.28% / 4 = 2.07%

r = Six months interest rate = (1+2.07%)^2 - 1 = 1.04182849 - 1 = 0.04182849 = 4.1828%

P = Semi annual payment = $2,950

PV = Debt value = $32,000

Payment = [r * PV] / [1 - (1+r)^-n]

$2,950 = [4.1828% * $32,000] / [1 - (1+4.1828%)^-n]

$2,950 = $1,338.496 / [1 - (1.041828)^-n]

1 - (1.041828)^-n = 0.453727458

(1.041828)^-n = 0.546272542

(1.041828)^n = 1.83058807

n = 14.754

Therefore, 15 payments needed to retire tha debt

Question b:

Cost of Debt for first five periods = (Total interest paid ) / Total debt

= ($1,338.496 + $1,271.090 + $1,200.864 + 1,127.702 + $1,051.479) / $32,000

= $5,989.63 / $32,000

= 0.187176

Question c:

Interest paid in 10th period is $619.79

Question d:

Amortization Schedule:
Period Opening Balance Installment payment Interest paid Principal amount repaid Closing Balance D=B*4.1828% E = C-D FEB-E 32